# Continuous emission ultrasound: a new paradigm to ultrafast ultrasound imaging

### v1

@@ 4.3 Potential applications
Regarding the potential improvement CEUI approach can provide for diverse ultrasound applications, to extend the 1D imaging scheme to a 2D imaging system using a multi-array probe would be a priority to investigate. For this purpose, the development of, a set of quasi-orthogonal emission waveforms with a beamforming method which will be capable to unmix and decode the contribution of each portion of each continuous emission on the recorded signals, will be investigated.
@@ 5 Conclusion
Future studies will be carried out to, first, extend the current framework to a highly temporally resolved two and three-dimensional monitoring of dynamic medium. This will raise some challenges both for the generation of adequat and optimized continuous waveforms and for the reconstruction step. In parallel, in vitro and in vivo experiments protocols will be developed to assess the feasibility of the method and to evaluate its performances on real data.

### v2

@@ 4.3 Potential applications
Regarding the potential improvement CEUI approach can provide for diverse ultrasound applications, to extend the 1D imaging scheme to a 2D imaging system using a multi-array probe would be a priority to investigate. An implementation based on the SISO CEUI framework combined with an approach similar to a full-matrix STA, a pulse compression and a temporal compounding, may be intended. For this purpose, the development of, a set of quasi-orthogonal emission waveforms with a beamforming method which will be capable to unmix and decode the contribution of each portion of each continuous emission on each recorded signal, will be investigated.
@@ 5 Conclusion
Future studies will be carried out to, first, extend the current framework to a highly temporally resolved two and three-dimensional monitoring of dynamic medium. This will raise some challenges both for the generation of adequate and optimized continuous waveforms and for the reconstruction step and will rely on the signal model in Appendix I. In parallel, in vitro and in vivo experiments protocols will be developed to assess the feasibility of the method and to evaluate its performances on real data.